OpenStudy (anonymous):

Solve x + 3y = 9 3x - 3y = -13 (-1,10/3 ) (1, -10/3 ) (-1,3/10 ) (1,-3/10 ) HELPPP PLEASEE

jigglypuff314 (jigglypuff314):

system of equations: elimination method x + 3y = 9 3x - 3y = -13 add as columns ------------ so 4x + 0 = -4 4x = -4 solve for x

jigglypuff314 (jigglypuff314):

then once you get x plug it into x + 3y = 9 to solve for y :)